# Learning Noise-Robust Stable Koopman Operator for Control with Physics-Informed Observables

### v2

@@ Abstract
We propose a novel learning framework for Koopman operator of nonlinear dynamical systems that is informed by the governing equation and guarantees long-time stability and robustness to noise. In contrast to existing frameworks where either ad-hoc observables or blackbox neural networks are used to construct observables in the extended dynamic mode decomposition (EDMD), our observables are informed by governing equations via Polyflow. To improve the noise robustness and guarantee long-term stability, we designed a stable parameterization of the Koopman operator together with a progressive learning strategy for roll-out recurrent loss. To further improve model performance in the phase space, a simple iterative strategy of data augmentation was developed. Numerical experiments of prediction and control of classic nonlinear systems with ablation study showed the effectiveness of the proposed techniques over several state-of-the-art practices.
@@ I Introduction
Data-driven methods have shown promising potential in modeling and control of nonlinear systems [1, 2]. The Koopman operator, as initially proposed by Bernard Koopman [3], plays a pivotal role by providing a global linear representation of nonlinear dynamical systems [4, 5]. Interested readers are encouraged to read several excellent reviews of Koopman operator [1, 6, 7, 8, 9]. The Koopman operator has been applied across diverse fields, including robust COVID-19 prediction [10], reinforcement learning [11], soft robotics control [12], stability analysis of genetic toggle switches [13], and modeling turbulent shear flows [14].
@@
The biggest benefit of Koopman operator for dynamics and control community is to facilitate the use of well-established linear control/observer techniques for the nonlinear systems [15, 16, 17], provided that a good approximation of global linear representation from data is learned. For example, model predictive control (MPC) is typically computationally expensive due to the need to solve non-convex optimization online. Traditional linearization techniques that approximate the nonlinear system around a local reference point might result in suboptimal MPC performance and could struggle with significant deviations from this point. On the other hand, the linear representation from Koopman operator can be used to transform the non-convex optimization of the nonlinear MPC into convex optimization of a linear MPC as shown by [18, 19]. In addition to MPC, such linear representation from Koopman operator can also help LQR [16], anomaly detection [20] and fault tolerant control [21] to nonlinear systems.
@@
Despite the several successful application of Koopman operator, there still remain several challenges. The first challenge is to find an appropriate invariant subspace for the Koopman operator. This involves selecting suitable basis functions that span an invariant subspace. Common choices of basis functions include radial basis functions (RBFs) [22] or polynomial functions [23]. Interestingly, Otto et al. [24] chose to learn the latent linear model and hidden state jointly, thereby eliminating the need to explicitly specify a dictionary of functions, as required by existing methods like Extended Dynamic Mode Decomposition (EDMD). However, one has to perform a non-trivial optimization for the linear dynamics and hidden states, which are unique for each single trajectory.
@@
Recently, the concept of Polyflow was introduced by [25]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept, which we refer to as physics-informed observables, is closely related to the derivative-based approximation of Koopman operator [26] except closure is considered. Despite the success of initial attempt of Polyflow for MPC [27], simply minimizing the one-step ahead error will nevertheless be prone to noisy data. The second challenge is to ensure long-term prediction accuracy. Wu et al. [28] reveal that under mild noise and nonlinearity, DMD can struggle to accurately predict the dynamics. Although the use of roll-out techniques for recurrent loss has proven beneficial when data is limited and noisy both in the deep learning [29, 30, 31] and the reduced-order modeling community [32], optimizing the recurrent loss can be tricky and prone to poor local minimum in practice. The third challenge in Koopman operator-based control is to ensure the stability of the Koopman operator itself since the real part of any continuous-time eigenvalue of the Koopman operator of any physical system should not be positive [31, 33]. If the forward pass becomes unstable, the backward pass will be unstable as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling especially in the high-dimensional phase space and, consequently, poor control performance. Besides, it is important to note that we do not attempt to address other challenges such as the convergence of Koopman operator. Interested readers are referred to this excellent paper by Colbrook and Mezić [34].
@@
In order to address the above issues, the goal of this paper is to develop a noisy-robust learning framework for Koopman operator with physics-informed observables. Specifically, the contributions of this work are:
  • approximating Koopman invariant subspace with physics-informed observables from Polyflow,
@@
The structure of the paper is as follows: Section II describes the proposed framework, including the use of Polyflow basis and the stabilized formulation. Section III presents the results of our experiments, demonstrating the efficacy of our approach. Section IV discusses the outcomes of applying iterative data augmentation on Koopman based control tasks. In Section V, we conduct an ablation study to evaluate the contributions of the various methods used in this paper. Section VI provides the conclusion of the paper.
@@ II-A Lifted Linear Predictor
Consider a nonlinear discrete-time dynamical system:
@@
where $f(\cdot)$ represents a generally nonlinear mapping, and $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ denote the state and input of the system, respectively. The objective is to predict the trajectory of the system, given an initial condition $x_{0}$ at time $t_{0}$, and a sequence of inputs $\{u_{i}\}_{i=0}^{N_{p}}$, where $N_{p}$ is the prediction horizon length of interest. To achieve this, we define the lifted linear predictor as a linear dynamical system determined by a triplet of matrices $(A,B,C)$:
@@
$z_{k+1}$ $=Az_{k}+Bu_{k},$
$\hat{x}_{k}$ $=Cz_{k},$
where $z\in\mathbb{R}^{N_{\phi}}$ is the lifted state, $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{N_{\phi}}$ is the lifting mapping, and $\hat{x}$ is the prediction of the original system’s state $x$. The lifting mapping $\Phi(\cdot)$ is defined as:
$\Phi(x)=[\phi_{1}(x),\phi_{2}(x),\ldots,\phi_{N_{\phi}}(x)]^{\top}\in\mathbb{R}^{N_{\phi}},$ (3)
where $\phi_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a scalar function referred to as an observable.
It is important to note that under mild assumpesions, the predictor could also take a bilinear form:
$z_{k+1}$ $=Az_{k}+Bz_{k}u_{k},$ (4)
$\hat{x}_{k}$ $=Cz_{k},$
as noted in several works [35, 36, 37, 38]. While it is natural to extend our work to accommodate the bilinear form, we avoid this for simplicity and maintain readiness for application in linear control systems.
@@ II-B Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor via state-space lifting, we briefly outline the Koopman operator approach for analyzing controlled dynamical systems. We consider the following discrete-time nonlinear system with control:
$x^{+}$ $=f(x,u)$ (5)
where, $x\in\mathbb{R}^{n}$ is the state of the system and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the system state of the next time step. We define the extended state comprising of the state $x$ of system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
@@
$\chi^{+}=F(\chi):=\begin{bmatrix}f(x,u(0))\\ \mathcal{S}\mathbf{u}\end{bmatrix},$ (6)
where $\mathcal{S}$ is the left shift operator, i.e., $(\mathcal{S}\mathbf{u})(i)=u(i+1)$, and $u(i)$ denotes the $i$th element of the sequence $\mathbf{u}$. The Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ associated to eq. 6 is defined by
@@
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$. Despite Koopman operator $\mathcal{K}$ being a linear operator fully describing the time evolution of any observable function on the non-linear dynamical system, it is inherently infinite dimensional that makes learning Koopman operator challenging.
II-C Extended DMD for Control with Physics-Informed Observables
For practical purposes, we adapt to the Extended Dynamic Mode Decomposition (EDMD) [39] approach followed by Korda and Mezić [18] for control. The EDMD provides a straightforward method to approximate the Koopman operator from data, thereby obtaining the lifted linear predictor in eq. 2. Since our goal is not to predict the control sequence, the observable function will depend only on the system states. Thus, the process begins by selecting $N_{\phi}$ observables $\phi_{1},\ldots,\phi_{N_{\phi}}$ and forming the lifted state $z$ from the measured state $x$ of the system (1) as follows:
@@ II-C Extended DMD for Control with Physics-Informed Observables
The one-step prediction loss based EDMD algorithm assumes that a collection of data $(x_{j},u_{j},x_{j}^{+}),j=1,\ldots,K$ satisfying $x_{j}^{+}=f(x_{j},u_{j})$ is available and seeks matrices $A$ and $B$ minimizing
@@
To select suitable basis functions $\Phi$ that span an invariant subspace, we are motivated by Polyflow basis functions from Wang and Jungers [27]. When the dynamics $f(x,u)$ is known, we can use a special set of basis functions $\Phi(x)=T_{N}(x)$, known as Polyflow basis functions of order $N\in\mathbb{Z}^{+}$ which is defined,
$T_{N}(x)$ $:=\begin{bmatrix}x\\ f(x,0)\\ f(f(x,0),0)\\ \vdots\\ f^{(N-1)}(x,0)\end{bmatrix}.$ (9)
@@
Using these Polyflow basis functions to lift the system state, we can derive the following system, which aligns with equation eq. 2
$z_{0}$ $=T_{N}(x_{0})$ (10)
$z_{k+1}$ $=Az_{k}+Bu_{k},$ (11)
$\hat{x}_{k}$ $=CT_{N}(x).$ (12)
Note that the matrix $C$ is typically known from the selection of $T_{N}(x)$.
Despite the similarity, time delay embedding [40] and Polyflow offer slightly distinct approaches to state lifting. Time delay embedding relies on observing and recording the physical system’s state over time to embed the state vector. In contrast, Polyflow directly utilizes the governing equations of the system, enabling an instantaneous embedding of the state without accessing the history of the physical system. Our framework assumes a known governing equation, but for systems with unknown dynamics, methods such as SINDy [41] or Neural ODE [42] can be employed as a surrogate of the governing equation.
II-D Recurrent Roll-out Loss
To enhance long-term prediction accuracy, we implement a recurrent roll-out loss instead of using one-step prediction loss eq. 8. For this, we consider a set of $M$ trajectories, each consisting of $R$ steps. Let $x_{i,j}$ denote the state at the $i$-th step of the $j$-th trajectory, where $i=0,1,\ldots,R$ and $j=1,2,\ldots,M$. The following optimization problem is then formulated:
$\underset{A,B}{\min}\sum_{j=1}^{M}\sum_{r=1}^{R}\|T_{N}(x_{r,j})-A^{r}T_{N}(x_{0,j})-\sum_{k=0}^{r-1}A^{r-k-1}Bu_{k,j}\|_{2}^{2}$ (13)
where, $A$ and $B$ are the matrices to be determined, and $u_{i}^{j}$ denotes the input at the $i$-th timestep of the $j$-th trajectory. To determine an appropriate basis size $N$, we solve the optimization problem (13) iteratively, increasing $N$ until the approximation error is sufficiently small.
II-E Curriculum Learning

### v5

# Learning Noise-Robust Stable Koopman Operator for Control with Hankel DMD

@@ Abstract
We propose a noise-robust learning framework for the Koopman operator of nonlinear dynamical systems, with guaranteed long-term stability and improved model performance for better model-based predictive control tasks. Unlike some existing approaches that rely on ad hoc observables or black-box neural networks in extended dynamic mode decomposition (EDMD), our framework leverages observables generated by the system dynamics, when the system dynamics is known, through a Hankel matrix, which shares similarities with discrete Polyflow. When system dynamics is unknown, we approximate them with a neural network while maintaining structural similarities to discrete Polyflow. To enhance noise robustness and ensure long-term stability, we developed a stable parameterization of the Koopman operator, along with a progressive learning strategy for rollout loss. To further improve the performance of the model in the phase space, a simple iterative data augmentation strategy was developed. Numerical experiments of prediction and control of classic nonlinear systems with ablation study showed the effectiveness of the proposed techniques over several state-of-the-art practices.
@@ I Introduction
Data-driven methods have shown promising potential in modeling and control of non-linear systems [1, 2]. The Koopman operator, as originally proposed by Bernard Koopman [3], plays a pivotal role by providing a global linear representation of nonlinear dynamical systems [4, 5]. Interested readers may refer to several excellent reviews on the Koopman operator [1, 6, 7, 8, 9]. The Koopman operator has been applied in various fields, including robust prediction of COVID-19 [10], reinforcement learning [11], soft robotics control [12], stability analysis of genetic toggle switches [13], and modeling turbulent shear flows [14].
@@
The biggest advantage of the Koopman operator for the dynamics and control community is its ability to facilitate the use of well-established linear control/observer techniques for nonlinear systems [15, 16, 17], provided that a good approximation of the global linear representation is learned from the data. For example, model predictive control (MPC) is typically computationally expensive as it requires solving a non-convex optimization problem online. Traditional linearization techniques that approximate the nonlinear system around a local reference point might result in suboptimal MPC performance and could struggle with significant deviations from this point. On the other hand, the linear representation of the Koopman operator enables the transformation of the nonconvex optimization problem in nonlinear MPC into a convex optimization problem in linear MPC, as demonstrated by [18, 19]. Beyond MPC, the linear representation provided by the Koopman operator can also benefit LQR [16], anomaly detection [20], and fault-tolerant control [21] for nonlinear systems.
@@
Despite its successful applications, several challenges remain in utilizing the Koopman operator. The first challenge is to find an appropriate invariant subspace for the Koopman operator. This involves selecting suitable basis functions that span an invariant subspace. Common choices of basis functions include radial basis functions (RBFs) [22] or polynomial functions [23]. Interestingly, Otto et al. [24] proposed learning the latent linear model and hidden state jointly, thereby eliminating the need to explicitly specify a dictionary of functions, as required by existing methods such as extended dynamic mode decomposition (EDMD). However, this approach requires solving a nontrivial optimization problem for the linear dynamics and hidden states, which are unique to each trajectory.
@@
Recently, the concept of polyflow was introduced by [25]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept is closely related to the derivative-based approximation of Koopman operator [26] except closure is considered. In earlier work, it was observed by Mezić [27] that instead of relying on arbitrary sets of observables, one could use observables generated by the dynamics, such as time delays that form a Krylov subspace, to study the spectral properties of the Koopman operator. This approach has also been applied in the context of Hankel-DMD methods. Despite the initial success of Polyflow for MPC [28], merely minimizing the one-step-ahead error remains susceptible to noise in the data. The second challenge is to ensure long-term prediction accuracy. Wu et al. [29] revealed that even under mild noise and nonlinearity, DMD struggles to accurately predict system dynamics. Several recent approaches have been proposed to address the challenges posed by measurement noise in learning Koopman operators. Sinha et al. [30] developed an online robust Koopman framework that employs a min-max optimization strategy to mitigate the effect of noisy streaming data, while Hao et al. [31] proposed a deep Koopman learning method that incorporates bounded noise considerations by updating the observable functions during training. However, these methods either require restrictive assumptions on noise bounds or face challenges in ensuring long-term prediction stability, which motivates the need for our stable parameterization and rollout-based learning strategy. Although the use of roll-out techniques for recurrent loss has proven beneficial when data are limited and noisy both in the deep learning [32, 33, 34] and the reduced-order modeling community [35], optimizing recurrent loss can be tricky and prone to poor local minimum in practice. The third challenge in Koopman operator-based control is to ensure the stability of the Koopman operator itself, since the real part of any continuous-time eigenvalue of the Koopman operator of any physical system should not be positive [34, 36]. If the forward pass becomes unstable, it inevitably destabilizes the backward pass as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling especially in the high-dimensional phase space and, consequently, poor control performance. Besides, it is important to note that we do not attempt to address other challenges such as the convergence of Koopman operator. Interested readers are referred to this excellent paper by Colbrook and Mezić [37].
@@
In order to address the above issues, the goal of this paper is to develop a noise-robust learning framework for the Koopman operator. Specifically, the contributions of this work are:
@@
The structure of the paper is as follows: Section II describes the proposed framework, including the use of the polyflow basis and the stabilized formulation. Section III presents the results of our experiments, demonstrating the efficacy of our approach. Section IV discusses the results of applying iterative data augmentation on Koopman-based control tasks. In Section V, we performed an ablation study to evaluate the contributions of the various methods used in this paper. Section VI provides the conclusion of the paper.
@@ II-A Lifted Linear Predictor
Consider a nonlinear discrete-time dynamical system,
@@
where $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ represents a general nonlinear mapping, and $x_{k}=x(t_{k})\in\mathbb{R}^{n}$ and $u_{k}=u(t_{k})\in\mathbb{R}^{m}$ denote the system state and control input at time $t_{k}$, respectively. Given an initial condition $x_{0}$ at time $t_{0}$ and a sequence of control inputs $\{u_{k}\}_{k=0}^{N_{p}}$, where $N_{p}$ is the length of the prediction horizon, the objective is to estimate the sequence of future states $\{x_{k}\}_{k=1}^{N_{p}+1}$.
To approximate the evolution of this nonlinear system using a linear framework, following the approach of [15], we consider a lifted linear predictor, modeled as a linear dynamical system parameterized by a triplet of matrices $(A,B,C)$:
@@
$z_{k+1}$ $=Az_{k}+Bu_{k},$ $k=0,1,\dots,N_{p},$
$\hat{x}_{k}$ $=Cz_{k},$ $k=1,\dots,N_{p}+1.$
where $z_{k}\in\mathbb{R}^{N_{\phi}}$ is the lifted state in an extended space of dimension $N_{\phi}$, and $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{N_{\phi}}$ is a nonlinear lifting function that maps the original state space to a higher-dimensional feature space. The vector $\hat{x}_{k}\in\mathbb{R}^{n}$ represents the estimated state of the system at time $t_{k}$.
The lifting function $\Phi(\cdot)$ is defined as:
$\Phi(x)=\begin{bmatrix}\phi_{1}(x)\\ \phi_{2}(x)\\ \vdots\\ \phi_{N_{\phi}}(x)\end{bmatrix}\in\mathbb{R}^{N_{\phi}},$ (3)
where each function $\phi_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is an observable, capturing potentially nonlinear transformations of the state. The choice of observables $\{\phi_{i}(x)\}_{i=1}^{N_{\phi}}$ determines the accuracy of the linear approximation.
It is important to note that, under certain conditions (e.g., control-affine systems), one can also adopt a bilinear model realization [38, 39, 40, 41, 42]. Suppose there exist coefficient matrices $A\in\mathbb{R}^{N_{\phi}\times N_{\phi}},\quad B_{i}\in\mathbb{R}^{N_{\phi}\times N_{\phi}},\quad C\in\mathbb{R}^{n\times N_{\phi}},$ such that for $z_{k}=[z_{k}^{(1)},\dots,z_{k}^{(N_{\phi})}]^{\top},\quad x_{k}=[x_{k}^{(1)},\dots,x_{k}^{(n)}]^{\top},\quad u_{k}=[u_{k}^{(1)},\dots,u_{k}^{(m)}]^{\top},$ the dynamics takes the form
$z_{k+1}$ $=Az_{k}+\sum_{i=1}^{m}u_{k}^{(i)}B_{i}z_{k},$ (4)
$x_{k}$ $=C\,z_{k},$
then $\bigl{(}A,\{B_{i}\}_{i=1}^{m},C\bigr{)}$ is said to be a bilinear realization of the system. While extending our approach to the bilinear setting is natural, we opt here for a linear formulation for simplicity and seamless integration with standard control methods.
@@ II-B Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor via state-space lifting , we briefly outline the Koopman operator approach for analyzing a controlled dynamical system, following the approach of [15]. We consider the following discrete-time nonlinear system with control,
$x^{+}$ $=f(x,u),$ (5)
where, $x\in\mathbb{R}^{n}$ is the state of the system and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the state of the system for the next time step. We define the extended state comprising of the state $x$ of system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
@@
$\chi^{+}=F(\chi):=\begin{bmatrix}f(x,\mathbf{u}(0))\\ \mathcal{S}\mathbf{u}\end{bmatrix},$ (6)
where $\mathcal{S}$ is the left shift operator, i.e., $(\mathcal{S}\mathbf{u})(i)=\mathbf{u}(i+1)$, and $u(i)$ denotes the $i$th element of the sequence $\mathbf{u}$. The Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ associated with eq. 6 is defined by
@@
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$. Despite the fact that the Koopman operator $\mathcal{K}$ is a linear operator that fully describes the time evolution of any observable function in a non-linear dynamical system, it is inherently infinite-dimensional which makes learning the Koopman operator challenging.
II-C Extended DMD for Control with Dynamics-Driven Observables
For practical purposes, we adapt the Extended Dynamic Mode Decomposition (EDMD) [43] approach followed by Korda and Mezić [18] for control. The EDMD provides a straightforward method to approximate the Koopman operator from the data, thereby obtaining the linear predictor lifted in eq. 2. Since our goal is not to predict the control sequence, the observable function will depend only on the state of the system. Thus, the process begins by selecting $N_{\phi}$ observables $\phi_{1},\ldots,\phi_{N_{\phi}}$ and forming the lifted state $z$ from the measured state $x$ of the system (1) as follows:
@@ II-C Extended DMD for Control with Dynamics-Driven Observables
The one-step prediction loss based EDMD algorithm assumes that a collection of data $(x_{j},u_{j},x_{j}^{+}),j=1,2,\ldots,K$ satisfying $x_{j}^{+}=f(x_{j},u_{j})$ is available and seeks matrices $A$ and $B$ minimizing
@@
II-D Polyflow Basis Functions for EDMD
To construct suitable basis functions $\Phi$ that span an invariant subspace for the Koopman operator, we leverage Polyflow basis functions, inspired by the work of Wang and Jungers [28], as well as Krylov subspace methods discussed by Mezić [27]. These basis functions provide a physics-informed approach to lifting the system state, ensuring a structured representation of the non-linear dynamics.
II-D1 Known Dynamics
When the system dynamics $f(x,u)$ is explicitly known, we define the Polyflow basis functions of order $N$, denoted as $T_{N}(x)$, as follows:
$T_{N}(x):=\begin{bmatrix}x\\ f(x,0)\\ f(f(x,0),0)\\ \vdots\\ f^{(N-1)}(x,0)\end{bmatrix}\in\mathbb{R}^{N_{\phi}}.$ (9)
Here, $f^{(i)}(x,0)$ represents the $i$-th recursive application of $f$ with zero input, capturing the intrinsic system evolution. The dimension of the lifted state is $N_{\phi}=N\times n$.
@@ II-D1 Known Dynamics
Using these Polyflow basis functions to lift the system state, we obtain the following linear system representation, consistent with the lifted linear predictor in eq. 2:
$z_{0}$ $=T_{N}(x_{0}),$ (10)
$z_{k+1}$ $=Az_{k}+Bu_{k},$
$\hat{x}_{k}$ $=Cz_{k}.$
The matrix $C$ is typically determined by the selection of $T_{N}(x)$ and encodes the mapping from the lifted state space back to the original state space.
Connection to Time-Delay Embedding
While both time-delay embedding [44] and Polyflow lift the state into a higher-dimensional space, their methodologies differ. Time-delay embedding constructs an extended state vector by recording historical state observations over multiple timesteps, effectively capturing past system behavior. In contrast, Polyflow directly leverages the governing equations to generate an instantaneous lifted representation without requiring historical state trajectories. This fundamental distinction makes Polyflow particularly suitable for embedding systems where access to past observations is limited or infeasible.
II-D2 Unknown Dynamics
If the governing dynamics $f(x,u)$ is unknown, we approximate them using a neural network $g{{}_{\theta}}(x):\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ parameterized by $\theta$, while maintaining the recursive structure of Polyflow. The corresponding basis functions are then defined as:
$T_{N}(x):=\begin{bmatrix}x\\ g{{}_{\theta}}(x)\\ g_{\theta}(g{{}_{\theta}}(x))\\ \vdots\\ g_{\theta}^{(N-1)}(x)\end{bmatrix}\in\mathbb{R}^{N_{\phi}}.$ (11)
Here, $g_{\theta}(x)$ serves as a learned approximation of $f(x,0)$, recursively applied to construct the lifted state representation. The dimension of the lifted state is $N_{\phi}=N\times n$. This data-driven extension allows Polyflow to be applied even in scenarios where an explicit model of the system dynamics is unavailable.
II-E Roll-out Loss
To improve long-term prediction accuracy and noise resistance in training data, we employ a multi-step rollout loss instead of the one-step prediction loss eq. 8.
Consider a data set consisting of $M$ trajectories, each trajectories consisting of $R$ roll-out steps, $\{(x_{0,j},x_{1,j},\dots,x_{R,j},u_{0,j},\dots,u_{R-1,j})\}_{j=1}^{M}$, where $x_{i,j}$ denotes the state of the system at the $i$th time step of the $j$-th trajectory, and $u_{i,j}$ is the corresponding control input.
The optimization problem for learning the system matrices $A$ and $B$ is then formulated as:
$\underset{A,B}{\min}\sum_{j=1}^{M}\sum_{r=1}^{R}\left\|T_{N}(x_{r,j})-A^{r}T_{N}(x_{0,j})-\sum_{k=0}^{r-1}A^{r-k-1}Bu_{k,j}\right\|_{2}^{2},$ (12)
where $A$ and $B$ are the unknown system matrices to be learned, and $T_{N}(x)$ represents the Polyflow basis function transformation of the state.
The hyperparameters $N$, and $R$ are typically determined through hyperparameter tuning to balance model complexity and predictive performance. Adjusting these parameters is a trade-off: increasing $N$ can capture more complex dynamics but may lead to overfitting if not managed carefully, while higher $R$ values typically enhance long-term prediction accuracy at the cost of increased computational burden.
II-F Curriculum Learning